OpenStudy (anonymous):

n the manufacturing process of sulfuric acid, sulfur dioxide is reacted with oxygen to produce sulfur trioxide. Using the equation, 2SO2 (g) + O2 arrow 2SO3 (g), if 128 g of sulfur dioxide is given the opportunity to react with an excess of oxygen, but only produces 144 g of sulfur trioxide, what is the percent yield of this reaction?

OpenStudy (anonymous):

@Zeta

OpenStudy (anonymous):

I know you can use the limiting reagent to solve it...is sulfur dioxide the limiting reagent? If so....eh, I have no idea what the hell I'm doing :(

OpenStudy (anonymous):

If you have 128 grams of sulphur dioxide, number of moles of sulphur dioxide = 128/(32 + 32) = 2 moles Your reaction is \(\rm 2 SO_2 + O_2 \to 2SO_3\), so the expected yield on the right hand side is 2 moles. 2 moles of \(\rm SO_3\) = 2 * (molar mass of SO3) = 2 * (32 + 48) = 2 * 80 = 160 grams. So you would expect 160 grams of sulphur trioxide, right?

OpenStudy (anonymous):

Sulphur dioxide is the limiting reagent here, but we don't need to think about limiting reagents here.

OpenStudy (anonymous):

makes sense...mostly how did you get again? sorry, ust don't want to keep asking the same kinds of questions over and over

OpenStudy (anonymous):

get *32 sorry, lag